# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

### v4

@@ VI-A1 Automated Cross-Platform Integrations
To fully leverage the flexibility of next-gen tools, practitioners need integration modules for popular CI/CD systems, registries, and orchestration platforms. Missing integration and sparse documentation forces users to implement custom connectors, increasing the risk of misconfiguration and potential security vulnerabilities that undermine the very guarantees signing is meant to provide.
VI-A2 Need for Inter-Platform Standards
@@
Standardization of information exchange is a standard problem for new classes of software infrastructure e.g., , SBOM [88]. These problems often harm integration capabilities. Next-generation signing tools would gain from common exchange standards—for example, uniform schemas for transparency logs or identity assertions. Projects like Diverfiy [89], which harmonize identity provisioning across multiple OIDC providers, are a promising start. Similar efforts are needed to streamline data flows between signing components, transparency services, and external tooling ecosystems.
VI-A3 Verification Workflow
@@
A key security property in a software supply chain is the transparency of actors and artifacts, which requires verifying signer identities [4]. Previous studies have highlighted challenges in signature verification across both organizational and open-source contexts [39, 5]. Next-generation tools like Sigstore integrate multiple verification methods, such as the Cosign CLI and the transparency log, but most participants still reported usability issues with the transparency log, and documentation indicates a lack of automation for other verification methods. Further automating the verification workflow would enhance the security benefits of software signing with this class of tools.
VI-B Usability Patterns of Sigstore vs Non-Sigstore users
Our results reveal different drives in adoption motivations and shared usability concerns between Sigstore users and non-users. Sigstore adopters cited contributions to the Sigstore project (Human/Social factor) and macroenvironmental factors—alongside negative experiences with legacy tools—as their primary drivers. In contrast, non-users emphasized organizational considerations when choosing their signing solution.
Despite these differences, both groups reported similar technological usability issues, particularly around integration with other systems and infrastructure transparency. Notably, non-users frequently highlighted signer identification challenges, whereas Sigstore users did not—underscoring the clear usability advantage of next-generation tooling.
VI-C Privacy Concerns
Privacy concerns emerged as a big issue among users of Sigstore’s Rekor transparency log module, with 5 out of 6 participants who reported drawbacks citing privacy as a significant concern. Privacy concerns were also major issues influencing users not to consider sigstore as suitable for enterprise. This issue also affects the use of OIDC for keyless signing. While this challenge is not unique to Sigstore [71] and is a recurring issue in software signing and modern cryptography, it underscores the need for continued innovation to address privacy-preserving mechanisms that balance transparency with user privacy, ensuring broader adoption and trust in cryptographic tools like Sigstore. While works like Speranza[71] has been proposed, efforts should focus on integrating these solutions to tools like Sigstore.
VII Threats to Validity
@@
To discuss the limitations of our work, we follow Verdecchia et al.’s [90] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our experiences and perspectives may have influenced our research (positionality).
@@
External Validity: The primary bias to our work is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects, and feel that that result — a focused treatment of usability in Sigstore — is a valuable contribution to the literature. A second threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. We attempted a survey (details are in section -E) but received only a few responses, echoing the challenges of surveying specialized communities [91]. A third threat is our study’s temporal scope. Interviews were conducted between November 1, 2023, and February 9, 2024, capturing Sigstore users’ experiences during that period. Since then, Sigstore has introduced some changes, but not the core workflows and components. This suggests our findings remain relevant.
@@
VIII Conclusion
Software signing guarantees provenance via cryptographic primitives, reducing supply-chain attacks. Next-generation tools like Sigstore automate traditional challenges, yet their usability remains underexplored. We interviewed practitioners to examine Sigstore’s usability, identifying key adoption factors. Our findings show that Sigstore’s identity-based workflows improve usability, but integration complexities, privacy concerns, and organizational constraints hinder adoption. We recommend toolmakers provide official integration plugins and privacy controls for transparency logs. These insights inform automation improvements, promoting secure, usable next-generation signing in automated software engineering.
IX Acknowledgments
@@
B: Software Supply Chain
Failure Experienced
B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process? How were these addressed? (Alternatively – How did this affect the decision to implement software signing?
[…]
B-5: What is your team’s greatest source of software supply chain security risk? Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies? How does the presence of a signature influence this decision? How is the authenticity of this signature verified? Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects? Possible Follow-up – Can you describe these methods/practices?
@@
C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
@@
C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)? how does the team use software signing to protect its source code (what parts of the process is signing required e.g. Commit signing)?
Build Process C-6: (How) Does the team utilize signing in its build process?
Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [92]).
@@
D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
@@
D-2: What factors did the team consider before adopting this tooling over others?
@@
D-3: What was the team’s previous signing practice before the introduction of current tooling?
@@
D-4: How does your team implement this tooling (which components of tool does the team use)?
D-5: Have you encountered any challenge(s) using this tool of choice? How have you coped with this limitation(s)? Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
-C Codebooks
@@
Table VII describes our codebook, with code definition, and example quote tagged with that code.
TABLE VII: Excerpts from the final codebook we used for our analysis.
@@
We present a saturation analysis of our studies in  Figure 7. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 12. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trend line (unique number of codes).
-E Survey Attempt
@@
What signing tool did your team use prior to the introduction of Sigstore?: Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
@@
In which of the following ways have you implemented Sigstore?: Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
Do you use or require Sigstore (signing) at any of the following stages of your product?: Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared the "Build/Deployment/final product" stage as part of their answer.
@@
-F Sigstore Components Most Commonly Used by Our Sample Population
TABLE VIII: Sigstore Functionalities Used by Subjects.
@@
-G Expanded Demographics Table
TABLE IX: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight the type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products than the one stated.
@@
-H Factors Influencing Sigstore Adoption (Sigstore Users)
TABLE X: Reasons Practitioners Choose or Switch to Sigstore Before Adoption. Practitioners’ decisions are driven primarily by human/social factors—such as community contribution, frustration with legacy tools, and usability issues—supplemented by Sigstore’s technological capabilities and reinforced by macroenvironmental pressures like regulations and trust in the CNCF ecosystem.

### v5

@@ 6.3 Lessons for Next-Gen Signing Usability
Automated Cross-Platform Integrations: To fully leverage flexibility, practitioners need maintained integration modules for popular CI/CD systems, registries, and orchestration platforms. Missing integrations and sparse documentation force teams to build custom connectors, increasing misconfiguration risk and undermining security guarantees.
@@
Need for Inter-Platform Standards: Standardization of information exchange is a typical problem for new infrastructure (e.g., SBOM [101]). Next-generation signing would benefit from uniform schemas for transparency events and identity assertions. Recent efforts like DiVerify [57], which harmonizes identity provisioning across multiple OIDC providers, take a step toward smoother data flows between signing components, transparency services, and external toolchains.
@@
Verification Workflow: A core supply-chain property is transparency of actors and artifacts, which requires reliable, usable verification of signer identities [15]. Prior work reports verification pain in both organizational and open-source settings [42, 71]. Although Sigstore provides multiple verification paths (Cosign CLI, transparency log), participants reported log-related usability issues and limited automation elsewhere. Further automating verification would better realize the security benefits of this tool class.
Privacy Concerns: Privacy concerns emerged prominently around Sigstore’s Rekor transparency log: five of the six participants who reported concerns with Rekor, cited privacy as a significant issue. Similar concerns also surfaced for keyless signing via OIDC. These are not unique to Sigstore [55] but reflect broader tensions in cryptography between transparency and confidentiality. This underscores the need for privacy-preserving designs (e.g., selective disclosure, redaction regimes, witness policies) that balance auditability with confidentiality. In Creswell’s terms, these concerns span macroenvironmental (policy/regulation), organizational (data governance and risk posture), and human/social (developer risk perception) factors, reinforcing that privacy is not solely a technical hurdle but a cross-factor adoption barrier.
Population Scope & Future Populations: Our study primarily sampled experienced practitioners operating within organizations, which shapes both the opportunities and frictions we observed. Usability needs likely differ for other communities and skill levels—e.g., individual open-source maintainers, small/startup teams, and security novices. In light of the biases from our particular sample, we recommend further research that (1) uses organizational contexts and user personas as a primary selection criterion; and (2) considers distinct signing use cases for a more fine-grained understanding.
7 Threats to Validity
@@
To discuss the limitations of our work, we follow Verdecchia et al.’s [97] recommendation to reflect on the greatest threats to our research. In addition, we discuss how our perspectives may have influenced our research (positionality).
@@
External Validity: The primary generalizability threat is that we recruited from a population biased toward the use of Sigstore. We viewed this as necessary in order to recruit sufficient subjects. A second threat arises from the sample size of our work, which is due to the challenges of recruiting expert subjects. A third threat is our study’s temporal scope. Interviews were conducted between November 1, 2023, and February 9, 2024, capturing Sigstore users’ experiences during that period. Sigstore has changed since then, but not the core workflows and components. We believe our findings remain relevant.
@@
8 Conclusion
Software signing underpins supply chain security by guaranteeing provenance and integrity, yet its effectiveness depends on usable tooling. Next-generation systems like Sigstore reduce traditional burdens of key management and signer identification, but our study shows that adoption is shaped by both strengths and frictions. Identity-based workflows, ephemeral keys, and CI/CD compatibility ease use, while integration gaps, transparency log privacy concerns, and organizational constraints remain barriers—especially in large enterprises.
By surfacing these factors, we show that usability is central to supply-chain security. Without attention to adoption barriers, even the best technical mechanisms will fail to achieve their security potential. Our findings reveal uneven maturity across Sigstore’s components, and suggest lessons for other signing toolmakers, e.g., OpenPubKey and SignServer which share similar design patterns. We recommend that toolmakers prioritize official integrations and configurable privacy controls to improve usability and trust. By addressing these challenges, next-generation signing tools can expand the coverage of verifiable artifacts and attestations, shrinking supply chain attack surfaces to improve software security.
9 Acknowledgments
@@
Ethical Considerations
Our study investigates the usability of next-generation software signing tools, focusing on Sigstore as an exemplar. We conducted semi-structured interviews with 17 security practitioners across 13 organizations, analyzing their perspectives on usability strengths, weaknesses, and adoption dynamics.
Stakeholders & Potential Harms.
Following the guidance of Davis et al. [25], we identify the following stakeholders.
  • Direct stakeholders: (i) Interview participants, whose responses could unintentionally expose their organization’s practices; (ii) The organizations employing these participants, which may face reputational, regulatory, or security risks if weaknesses are revealed; (iii) The research team, responsible for balancing transparency and confidentiality; and (iv) Adversaries, who could misuse disclosed weaknesses in signing tools or deployment practices.
  • Indirect stakeholders: (i) Other software-producing organizations, who may be affected if the study highlights industry-wide shortcomings in signing tool used by such organization; (ii) Standards bodies, whose guidelines may be challenged; (iii) Software consumers, whose trust in signed artifacts could be affected; and (iv) The broader public, for whom systemic weaknesses in signing infrastructure may undermine confidence in software supply chains.
Mitigations Implemented.
Recruitment Policies:Data were collected through voluntary interviews, with written consent obtained prior to participation. Recruitment was conducted through established professional networks; no unsolicited contact or spamming was used. Participants were free to decline or withdraw at any stage, and recruitment avoided any form of coercion. All organizations were treated equally in data collection and analysis. At no point were sensitive keys, cryptographic material, or detailed system configurations collected.
Content safeguards: All transcripts were anonymized; neither participant names nor organizational identifiers appear in published results. Findings are reported in aggregate at the thematic level (e.g., categories of usability issues) rather than specific configurations or vulnerabilities.
Privacy & data minimization: Only de-identified transcripts were retained; analysis outputs emphasize themes across participants, not individual accounts. No identifiable deployment details are included in the publication.
Operational safety: We deliberately excluded or generalized any descriptions of signing deployments that could aid adversaries, including specific weaknesses in Sigstore’s components. Where risks were mentioned (e.g., transparency log privacy concerns, private instance deployment issues), results are presented at the level of usability factors rather than detailed attack vectors.
Participant Welfare.
As mentioned earlier, all participants gave informed consent. Potential distress was minimal, though participants might experience discomfort when reflecting on organizational challenges. To mitigate this, questions emphasized practices and perceptions rather than failures, and responses were anonymized. The study did not involve deception.
Oversight and Consent Context.
The academic institution’s IRB reviewed and approved the protocol. Prior to consenting, all participants were informed of the study’s aims and the confidentiality safeguards in place.
Decision to Proceed and Publish.
We judged that the expected benefits—independent evidence on the usability of next-generation signing tools, insights into adoption barriers, and guidance for improving software supply chain security—outweigh residual risks under the mitigations described above. To minimize adversarial risk, we publish only aggregated themes, not sensitive technical details, ensuring that our findings strengthen rather than weaken trust in signing infrastructures.
Open Science
We acknowledge that USENIX Security has an open science policy: that authors are expected to openly share their research artifacts by default. The research artifacts associated with this study are:
  • Raw transcripts of interviews
  • Anonymized transcripts of interviews
  • Interview protocol
  • Codebook
Things we have shared: The interview protocol is a crucial part of any interview study, since it allows for the critical review of a study design as well as its replication. Our full interview protocol is included in Appendix˜B. Since it is semi-structured, we include all of the questions asked of all subjects, as well as examples of the follow-up questions we asked. We also share the codebook, with codes, definitions, and example quotes that we coded for each code — see Appendix˜C.
Things we cannot share: For subject privacy reasons, and for IRB compliance, we cannot share the raw transcripts. We are also unwilling to share the anonymized transcripts of the interviews. Given the high organizational ranks of many of our subjects, and the small size of the subject pool resulting from our recruiting strategy, we believe there is a high risk of de-anonymization even of anonymized transcripts. Therefore, we cannot share the anonymized transcripts.
Artifact Repository: An anonymized artifact containing our interview protocol, codebook, etc. is available at: https://github.com/nextgenusability/next-gen-usabilty
Appendix A Traditional Signing Tools Workflow
The signature creation process begins when the maintainer’s artifact is ready for submission. The signer generates a key pair that provides them with a private key for signing their software artifact and a public key for a verifier to validate the artifact’s signature. Once the artifact is signed, it must be submitted to a package registry (such as PyPi, npm, or Maven) or the intended party. During this phase, the signer is responsible for making their public key available and discoverable so future users can verify the signer’s identity. In the verification phase, the verifier retrieves the signer’s public key, uses it to decrypt the signature file, and checks for equivalency with the software artifact’s hashed digest. We show a typical traditional package signing workflow in Figure˜6 compared to a next-generation workflow in Figure˜7.
Figure 6: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
Figure 7: Typical workflow for next-generation software signing. The component author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package.
Appendix B Interview Protocol
Table˜1 gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section B & C of the protocol for completeness, although they were not analyzed in this study.
Table 6: Interview Protocol. Sections A–C reflect demographics, signing use/motivation context, and signing tool usability. We analyze part C in this study and use the other themes as context.
Theme Description / Questions
A:
Demographics
We ask basic demographic questions to establish participant experience and context (role, years of experience, team size, and major products).
A-1: What best describes your role in your team? (Security engineer, Infrastructure, Software engineer, etc.)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: What are the team’s major software products/artifacts?
B:
Signing Use &
Context
We ask about the organizational context of signing: why and how signing is used, what motivates adoption, what infrastructure supports it, and how it fits into broader supply chain security practices. Focus is on general practices, not tooling specifics.
B-1: Why did the team choose to implement software signing?
B-2: Is software signing the team’s major strategy to secure its supply chain? Any complementary efforts?
B-3: Are these strategies influenced by regulations or standards?
B-4: What challenges did your team face while integrating signing into supply chain practices?
B-5: How do team members contribute to the implementation of signing? Roles/responsibilities?
B-6: Do you believe signing (if implemented correctly) is sufficient to secure a supply chain?
B-7: How does the presence of a signature influence dependency selection?
B-8: How is authenticity of a signature verified in your process?
@@ Appendix B Interview Protocol
B-9: What influences your team to discontinue a dependency? Does signature status play a role?
@@
B-10: How does your team manage third-party dependency security and vulnerabilities?
B-11: Is signing a requirement for developers and contributors? If so, in which parts of the workflow?
B-12: How is signing applied in your build process?
B-13: How is signing applied to artifacts and deployment (binaries, repositories, pipelines)?
C:
Signing Tool
Usability
We ask about the signing tools themselves: what tools are in use, factors influencing adoption, challenges, and considerations of alternatives. Each question includes a brief explanation.
@@
C-1: What software signing tool does the team use? (Understanding current tool choices and components in practice.)
@@
C-2: What factors did the team consider before adopting this tool over others? (Identifying decision-making drivers.)
@@
C-3: What was the team’s previous signing practice before the current tool? (Capturing historical evolution of practices.)
@@
C-4: How does the team implement this tool (which components are in use)? (Understanding deployment in workflows.)
C-5: What challenges or limitations has the team faced with this tool? How did you cope with these? (Surfacing usability pain points and coping strategies.)
C-6: Has the team considered switching tools? If yes, which alternatives? (Exploring sustainability of current tool choices.)
Appendix C Codebooks
@@
Table˜7 describes our codebook, with code definition, and example quote tagged with that code.
Table 7: Excerpts from the final codebook we used for our analysis.
@@ Appendix D Code Saturation
We present a saturation analysis of our studies in  Figure˜8. The green trendline (saturation curve) was measured by identifying the cumulative unique codes present in each interview. We observe saturation at the eleventh interview, indicating that no new unique codes were identified after participant 12. Additionally, the blue trendline (number of codes) shows that the lowest number of codes was obtained from subject S4. This is also reflected in the orange trend line (unique number of codes).
Appendix E Survey Attempt
@@
What signing tool did your team use prior to the introduction of Sigstore? Among 7 participants who utilize Sigstore, 3/7 responded that they use PGP/GPG-based signing and 4/7 did not use any toolings prior to using Sigstore.
@@
In which of the following ways have you implemented Sigstore? Among 7 participants who utilize Sigstore, 3 responded to this question, with 2 responding that they use a private deployment of Sigstore and 1 responding that they use a public instance of Sigstore.
Do you use or require Sigstore (signing) at any of the following stages of your product? Among 7 participants who utilize Sigstore, 3 responded to this question. All 3 participants shared the "Build/Deployment/final product" stage as part of their answer.
@@
E.1 Sigstore Components Most Commonly Used by Our Sample Population
Table 8: Sigstore Functionalities Used by Subjects.
@@ E.1 Sigstore Components Most Commonly Used by Our Sample Population
E.2 Other Results
In this appendix we prresent a continuation of the results presented and shown in Figure˜4 to show the Organizational and usage context report from participants in medium-scale organizations.
Figure 9: Heatmap showing the organizational usability context for subjects from medium scale organizations (4 subjects). Similar to the one showed for small and large-scale organizations in Figure˜4. In comparing to Figure˜4, note that some columns are omitted in this figure because no subjects mentioned those aspects.
E.3 Expanded Demographics Table
Table 9: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives; “Technical leader” to senior, lead, partner, and principal engineers; and “Engineer” and “Manager” to more junior staff. We highlight the type of software subject’s immediate team produces. * refers to cases where subject’s team built other types of software in addition to the main stated. O refers to cases where subject belonged to multiple teams with other products than the one stated.
@@ E.3 Expanded Demographics Table
E.3.1 Organizational Demographics
We summarize additional organizational demographic information of our subjects in Table˜10. This table provides further context for the subject’s organizational size, type of software product, and how many of our subjects belonged to each organization.
Table 10: Organizational Demographics. To enhance anonymity, We only highlight the number of subjects in each organizational category. letters A-L is used to depict each organization.
Type Breakdown (#Organizations|#Subjects)
Organizational Size (Employee Size) Small (<100) (4/6), Medium (<1500) (3/4), Large (>1500) (6/7)
Product Area Digital technology (1/3), SSC Security (2/4), Social technology (1/1), Dev tools (1/1), Telecommunications (1/1), Cloud security (2/3), Aerospace Security (1/1), Internet services (2/2), Cloud + OSS Security (1/1), Cloud + Dev toools (1/1)
Subject Distribution A (2), B (3), C (1), D (1), E (2), F (1), G (1), H (1), I (1), J (1), J (1), K (1), L (1), M(1)
E.4 Factors Influencing Sigstore Adoption (Sigstore Users)
In this appendix, we present a tabular organization of the reasons given by Sigstore users for their teams’ and organizations’ adoption of Sigstore, structured according to Cresswell et al. ’s four-factor framework.
Table 11: Reasons Practitioners Choose or Switch to Sigstore Before Adoption. Practitioners’ decisions are driven primarily by human/social factors—such as community contribution, frustration with legacy tools, and usability issues—supplemented by Sigstore’s technological capabilities and reinforced by macroenvironmental pressures like regulations and trust in the CNCF ecosystem.